if every fourth trick- or- treater is a vampire, and every 6th a robot , how many of each will you see if 75 trick- or- treaters come to your door?
i need the steps on how u solve this problem.
THANK YOU:)

Found 2 solutions by edjones, josmiceli:
Answer by edjones(8007)

What steps?
75/4=18.75
75/6=12.5
So, 18 vampires and 12 robots will appear.
Ed

Answer by josmiceli(19441)

I just drew up the whole thing. 75 events can be shown as 15 groups
of 5 each (75 = 15*5). I wanted to see if the robots and vampires
would try to come at the same time
-------------------------------------------
1 2 3 4 5 6 7 8 9 0 I 2 3 4 5
V 1 1 R V 1 1 1 V R 1 1 V 1 1
2 2 2 V R 2 2 V 2 2 R V 2 2 2
3 3 V 3 3 R V 3 3 3 V R 3 3 V
R V 4 4 4 V R 4 4 V 4 4 R V 4
V R 5 5 V 5 5 R V 5 5 5 V R 5
-------------------------------------------
There are 15 groups of 5 each. Every 4th is a Vampire. Every 6th is
a Robot. If you look at the patterns of R's and V's they arediagonals
going in opposite directions, so a V & R can never end up in the same
position. Counting V's I get 19. Counting R's I get 12